OpenStudy (anonymous):

calculate the area of a trapezium ABFE. AB is 90cm , E F is 40 cm (height)

OpenStudy (anonymous):



OpenStudy (anonymous):

please help :(

OpenStudy (anonymous):

 area of a trapezium = 0.5*height(sum of parralel sides) =0.5*40(90+90) =20(180) =3600